OpenStudy (anonymous):

Factoring 4x²+8x+3=0

OpenStudy (anonymous):

(2 x+1) (2 x+3) = 0

OpenStudy (anonymous):

(2x + 1) (2x+3)=0 x=-1/2, -3/2

OpenStudy (anonymous):

4x^2 + 8x + 3 = 0 4x^2 + 2x + 6x + 3 = 0 2x(2x + 1) + 3(2x + 1) = 0 (2x + 3)(2x + 1) = 0 2x + 3 = 0 x = -3/2 2x + 1 = 0 x = -1/2